RUN 2B L1CAL MEETING MINUTES 24 January, 2006 Present o Fermilab L.Bagby, S.Cihangir, M.Cwiok, G.Ginther, S.Lammers, M.Mulhearn, V.O'Dell, A.Stone, R.Unalan, N.Varelas, T.Zmuda o FSU T.Adams o Indiana H.Evans o MSU M.Abolins, D.Edmunds, P.Laurens o Northeastern E.Barberis, A.Roe, D.Wood o York S.Beale (apologies to those I couldn't see) Welcome to (potential) New Member --------------------------------- o Selcuk Cihangir from CMS silicon Cable Testing Report: M. Mulhearn --------------------- o Testing of full installation of ADF to TAB cables - 1st step: testing of 80 "central" cables (of 240 total) . 4 bent pins: a Fermilab expert (Joshua) will help us straightening these probability of success is fairly hight . most of the rest are bad cables - all 240 cables . 30 problem cables . intermittent parity errors happening on ~1/2 of the cables in the system > problem appears to be at cable-ATC-ADF end (probably bad cables again) - Dan can test "flaky cables" at MSU . will send him an entire bundle o Note: need to start partitioning problems so that we can continue testing the rest of the system o Possible causes for these problem - cables appear to be very fragile - but have withstood >1 year of abuse at Nevis, so problems are probably caused by a specific type of mistreatment o More cables have been ordered - turnaround is ~6 weeks o Should start up daily integration meetings again - 9:15 am in the 1st floor kitchen area PREPARATIONS FOR ADM ON FEB. 8 ------------------------------ Hardware Integration: M. Mulhearn --------------------- o Cabling - see above o GAB to L2/L3 tests - L2: probably format changes - L3: data length appears to be changing event-by-event o 48V supply overcurrent protection o VME/SCL --> SCL communication: minor test Firmware Status: M. Mulhearn ---------------- o MSU: no changes needed o COOR to TAB/GAB code o New EM algorithm has been checked - more detailed tests needed o TAB to GAB data transmission glitches o GAB: need to add a subset of triggers for Feb.8 Online/COOR: P.Laurens ------------ see text linked to agenda for a full list of these items Monitoring: D.Wood ----------- see text linked to agenda for a full list Examine: S.Beale -------- see slides linked to agenda o online tests should be possible by Feb.8 note: the online part is unchanged from old system Pulser: P.Renkel ------- o no report Offline/Unpacker: T.Adams ----------------- from Todd... All of the major functionality of the unpacker code is written. The code still needs to be fully tested/certified on data containing multiple TABs and containing GAB data. This is ongoing. All that is needed to be ready for stable operations is good data to test it and a few hours of work. Beyond Feb. 8 we will want to find other places the unpacker will be used (L3, reco, etc) and make sure everything is provided. o will check what needs to be done for L3, reco, etc. Simulation of Algorithms: S.Lammers ------------------------- o see text linked off agenda for full list Estimated Rates: E.Barberis ---------------- From Ela... Rates studies are now done in the context of the v15 triggerlist working group, which meets twice a week, updating a trigger list which is now aiming at 300E30. Rates and thresholds studies are now, rightly, conducted at a larger level than the Run2b expert pool, and the tools used are the trigger rate tool, CAF trees with 2b objects, trigsim. I believe Marco has taken a very recent enhanced bias run. o question to think about: how much do we gain with the new system? (given that there will still be risks associated with the change) - need to develop a strategy for dealing with these types of questions - note: this will also be covered by the v15 ADM talk Installation Plans: L.Bagby/D.Edmunds ------------------- From Lyn... a) We had a meeting last week and reviewed an itemized task list from John and Alan for MCH1. Action items included the assignment of technician names from EE support to specific tasks and reordering a few tasks. Tasks without EE support will need L1_Cal group support. Once the updated task list from John is submitted, Lyn will fold in anything on Alan's list not already covered. Currently, Lyn is waiting on John's input. b) Based on Mike's bad cable result, new cables need to be integrated into the current cable bundles or hand laid. c) CRATER has been updated and needs to be launched and tested. d) The LV monitoring GUI needs to be launched. SES limits need to be defined. Guidance for SES alarms need written. e) Instructions for shifters need to be written. f) A special run needs to be done to confirm that reco doesn't crash. (This item is based on a conversation with Bill Lee and may have already been done??) Shutdown Start Date ------------------- from George... o new official date: Monday, Feb. 27 - unless we have another accelerator problem - scheduled to end June 2 . unless changed by zebra mussels (depends on the weather)